OpenStudy (anonymous):

du/dx y-4(x^2)+2=0 I got (8x/y) but that isnt accepted.....

hartnn (hartnn):

y-4(x^2)+2=0 dy/dx - 8x +0 = 0 dy/dx = 8x its just 8x

OpenStudy (anonymous):

so if y has no coefficient is just turns into 0?

hartnn (hartnn):

it didn't turn to 0, y turned to dy/dx after differentiating

hartnn (hartnn):

just to make sure, u want dy/dx in y-4(x^2)+2=0

hartnn (hartnn):

?

OpenStudy (anonymous):

well it turned into yy' right and the dervivative of y is 1 so is that what happened? and yes, thank you for your help

hartnn (hartnn):

the derivative of y isn't yy' its just y'

OpenStudy (anonymous):

okay i guess i just understood it wrong..... i have another question if you dont mind sqrt(x)+sqrt(y)=7 and y(9)=16 find y'(9) by implicit differentiation

hartnn (hartnn):

know implicit diff. ?

OpenStudy (anonymous):

yes, find y' as a function of x i came up with (-1/2x^(-1/2))/((1/2)y^(-1/2))

hartnn (hartnn):

wait, was that y' ?

OpenStudy (anonymous):

yes thats what i hope y' is

hartnn (hartnn):

yes, thats correct, y'=(-x^(-1/2))/(y^(-1/2)) now put ,x=9, y=16

OpenStudy (anonymous):

okay thanks i was confused as to what to plug in for y..... thanks again!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

if you wouldnt mind i have another question xy^12=1 at (1,-1) find the slope of the tangent line at this point. so i got x*12y^11 and tried to plug in the points but to no avail..... any suggestions?

hartnn (hartnn):

\(xy^{12}=1\) your derivative is incorrect, u need to use product rule.

hartnn (hartnn):

12xy^11y' + y^12 = 0 now put x=1,y=-1

OpenStudy (anonymous):

awesome, thanks!

hartnn (hartnn):

-12y'+1=0 y'=-1/12 u got this ?

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

to Find dy/dx in terms of x and y if 7 x y + 4 x + y = 12 would the next step be 7yy'+4+yy' ?

hartnn (hartnn):

for 7xy, it will be 7yy'+7y other terms are correct

OpenStudy (anonymous):

okay so there will be no x in the answer? and i am confused as to what to do once i have 7yy'+7y+4+yy'

hartnn (hartnn):

7yy'+7y+4+y' = 0 now isolate y' Note, last term was also incorrect, its just y'

OpenStudy (anonymous):

would it turn into y'=-4/7y or 2y'=-4/7y or neither?